roblem Consider a one period (discount) bond that pays $100 at the end of the year. 1. Ifthe current interest rate is 5 percent per-annum, what is the current price ofthe bond? Answer: Suppose the current interest rate was to fall to 2 percent, what is the effect on the price of the bond? 2. Answer: Problem 2
Answer #1

(1)

Current bond price ($) = Price after 1 year / (1 + Interest rate) = 100 / (1 + 0.05) = 100 / 1.05 = 95.24

(2)

When interest rate is 2%,

Current bond price ($) = Price after 1 year / (1 + Interest rate) = 100 / (1 + 0.02) = 100 / 1.02 = 98.04

Therefore, bond price will be higher by $(98.04 - 95.24) = $2.8.